In a family, a couple has a son and a
daughter. The age of the father is four times that
of his daughter and the age of the son is half of
that of his mother. The wife is 8 years younger to
her husband and the brother is seven years older
than his sister. What is the age of the mother?

Answer by stanbon(75887)

The age of the father is four times that
of his daughter and the age of the son is half of
that of his mother. The wife is 8 years younger to
her husband and the brother is seven years older
than his sister. What is the age of the mother?
-----
Equations:
f = 4d
s = (1/2)m
f = m + 8
s = d + 7
---------------------
Solve for "m"::::
---------------------
m = 2s
m = f -8
--------
2s = f-8 = 4d-8 = 4(s-7)-8
-----
So, 2s = 4s-28-8
2s = 4s - 36
2s = 36
s = 18 (son's age)
---
Since s = (1/2)m, mom's age = 36
-------------------------------------
Cheers,
Stan H.
----------------